OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

x-y=10--eqn 1 2x + y = 2--eqn2 eqn 1 + eqn 2=>3x=12 x=4 eqn 1=>4-y=10 y=4-10 =-6 Elimination method. Since we had y with both + and - signs and the cofficients were equal I thought it wud b fyn to use em.